OpenStudy (anonymous):

Please Help!!!! Verify the identity. cot(x=pi/2)=-tanx

OpenStudy (anonymous):

is this \(\cot(x-\frac{\pi}{2})=-\tan(x)\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you are allowed to use the "cofunction" identity you get it in one step

OpenStudy (anonymous):

otherwise, use the "subtraction angle" formula for cotangent

OpenStudy (anonymous):

the question did not specify what I can and cannot use to sole the problem

OpenStudy (anonymous):

one of the "cofunction" identities gives \[\cot(\frac{\pi}{2}-x)=\tan(x)\] since cotangent is odd, this tell you \[\cot(x-\frac{\pi}{2})=-\cos(\frac{\pi}{2}-x)=-\tan(x)\]

OpenStudy (anonymous):

elsewise you can use \[\cot(x-\frac{\pi}{2})=\frac{\cos(x-\frac{\pi}{2})}{\sin(x-\frac{\pi}{2})}\]

OpenStudy (anonymous):

\[Cot(A-B) = \frac{ cotA.CotB+1 }{ CotB-CotA }\]

OpenStudy (anonymous):

then use the "subtraction angle" formula top and bottom it is a pain to write but it will work

OpenStudy (anonymous):

or you can use what @Yahoo! wrote above

OpenStudy (anonymous):

could you explain the cofunction identity alittle more to me please?

OpenStudy (anonymous):

or you can use a right triangle 

OpenStudy (anonymous):

from the picture above, it should be clear more or less that \(\tan(x)=\cot(\frac{\pi}{2}-x)\)

OpenStudy (anonymous):



OpenStudy (anonymous):

\(\tan(x)=\frac{b}{a}=\cot(\frac{\pi}{2}-x)\)

OpenStudy (anonymous):

now since cotangent is odd, and since \(x-\frac{\pi}{2}=-(\frac{\pi}{2}-x)\) you have \[\cot(x-\frac{\pi}{2})=-\cos(\frac{\pi}{2}-x)=-\tan(x)\]

OpenStudy (anonymous):

ohh okay thank you so much!!!

OpenStudy (anonymous):

cofunction identities are \[\sin(\frac{\pi}{2}-x)=\cos(x)\] etc

OpenStudy (anonymous):

works with any "co"

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so as my final answer i would just write cot(x-pi/2)=-tan x cot(x-pi/2)= -cos(pi/2-x)=-tan x ?

OpenStudy (anonymous):

well we know that \[\cot(\frac{ \pi }{ 2 }-u)=\tan u\] So, we know that \[\cot(x-\frac{ \pi }{ 2 }) = -\tan x\] :o)